OpenStudy (anonymous):

The range of f(x) = cos(x) is y ≤ 0 Answer True False

OpenStudy (mathmate):

The given statement is "The range of f(x)=cos(s) is less than or equal to zero" Assuming y=f(x), then y is in the range of cos(x). Using your calculator, try values of cos(0), cos(10 degrees), etc. to see if the values that you get satisfies y\(\le\) 0. If \(any\) value you obtain does not satisfy the condition, it constitutes a counter example and the given statement is false.